OpenStudy (anonymous):

8x-1=4x + 23 what does x equal? please help!

OpenStudy (anonymous):

8x-1=4x+23 +1 +1 8x = 4x + 24 -4x -4x 4x=24 4x/4 = 24/4 x=6

OpenStudy (anonymous):

Do u get it??

OpenStudy (anonymous):

@foreverandalways

OpenStudy (anonymous):

i kinda get it, but im realy bad at math so its still confusing. but thankyou for your help! (:

OpenStudy (anonymous):